OpenStudy (anonymous):

Given the theorem: 1 + 3 + 5 + ... + (2n - 1) = n^(2) Show that it holds for n = 5.

OpenStudy (hoblos):

for n=5 => 2n-1 = 9 1+3+5+7+9=25=5^2

OpenStudy (anonymous):

ahh I see. thanks